OpenStudy (anonymous):

How would you solve this: Use those steps to find the equation of the line through the points (2, 5) and (-3, 1).

OpenStudy (anonymous):

step 1, find slope step 2, use point slope formula step 3, convert to slope intercept formula or standard formula

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

so would the answer for this be m=6 b=-4

OpenStudy (anonymous):

slope is incorrect m= 5-1/2-(-3) m=4/5

OpenStudy (anonymous):

how did you get that could you show me so I will understand

OpenStudy (anonymous):

look at the slope equation i gave you

OpenStudy (anonymous):

(2, 5) = (x2,y2) (-3, 1)=(x1,y1) x2=2, y2=5, x1=-3, y1=1

OpenStudy (anonymous):

1-5/-3-2= -4/-5 < --- i believe thats your slope of y=mxtb .. so y=-4/-5x + b .. you can check by using one of the coordinates

OpenStudy (anonymous):

-4/-5= 4/5